OpenStudy (anonymous):

Graph the first six terms of a sequence where a1 = 3 and d = -10.

OpenStudy (jdoe0001):

\(\large a{\color{brown}{ n}}=a_1+({\color{brown}{ n}}-1)d\qquad \begin{array}{llll} term&value \\\hline\\ a_{\color{brown}{ 1}}&-3\\ a_{\color{brown}{ 2}}&a_1+({\color{brown}{ 2}}-1)\cdot -10\\ a_{\color{brown}{ 3}}&a_1+({\color{brown}{ 3}}-1)\cdot -10\\ a_{\color{brown}{ 4}}&a_1+({\color{brown}{ 4}}-1)\cdot -10\\ a_{\color{brown}{ 5}}&a_1+({\color{brown}{ 5}}-1)\cdot -10\\ a_{\color{brown}{ 6}}&a_1+({\color{brown}{ 6}}-1)\cdot -10 \end{array}\)

OpenStudy (jdoe0001):

hmmm well \(\large a_{\color{brown}{ n}}=a_1+({\color{brown}{ n}}-1)d\qquad \begin{array}{llll} term&value \\\hline\\ a_{\color{brown}{ 1}}&-3\\ a_{\color{brown}{ 2}}&a_1+({\color{brown}{ 2}}-1)\cdot -10\\ a_{\color{brown}{ 3}}&a_1+({\color{brown}{ 3}}-1)\cdot -10\\ a_{\color{brown}{ 4}}&a_1+({\color{brown}{ 4}}-1)\cdot -10\\ a_{\color{brown}{ 5}}&a_1+({\color{brown}{ 5}}-1)\cdot -10\\ a_{\color{brown}{ 6}}&a_1+({\color{brown}{ 6}}-1)\cdot -10 \end{array}\)